# Annamaria Porreca Fabrizio Maturo Viviana Ventre

### v1

@@ 6.3.1 Fuzzy in-closeness centrality
where $\widetilde{A}_{ij}$ is the set of the best fuzzy edges to connect $i$ and $j$ ranked in decreasing order, s indicates the maximum number of steps considered evaluating the best path, and $w_{j}$ a vector of weights of the FOWA operator. Clearly, being a weighted average of fuzzy numbers, $\widetilde{M}_{v_{i},s}$ is also a fuzzy number.
Hence, we can define the fuzzy out-closeness centrality index as follows:
@@
Thus, differently from Equation 32 that is a relative frequency, $\widetilde{C}_{v_{i},s}$ is a centrality index that incorporates the fuzziness of the edges and thus remains fuzzy.
Another possible approach to define fuzzy in-closeness centrality, in a way that is more close to the classical index, would be to consider:
$\widetilde{C^{(2)}}^{in}_{v_{i},s}=\frac{1}{\widetilde{M}_{v_{i},s}}.$ (35)
However, in the following we refer to Equation 34 because our basic idea is to look for an immediate representation of the in-closeness concept rather than to identically reproduce the classical indexes using fuzzy numbers. In fact, Equation 34 is quite interpretable, because provides a measure of how a node is central in considering the fuzzy incoming edges. However, Equation 34, differently from 26 takes into account the concept of the best path and maximum number of steps to reach another node.
@@ 6.3.2 Fuzzy out-closeness centrality
Following the same reasoning, we can define the fuzzy in-closeness centrality index of the node $v_{i}$ as follows:
@@
where $\widetilde{Z}_{v_{i},s}$ is the average (fuzzy number) of the best paths from nodes $j$ to $j$ in decreasing order. Thus,
@@
is computed by applying the FOWA operator to fuzzy numbers representing the edges from the generic nodes $j$ and the node $i$, which we consider centrality.
As before, the vector of weights can be of different types. From now on, we will compute the average fuzzy number $\widetilde{M}_{v_{i},s}$ or $\widetilde{Z}_{v_{i},s}$ using $w=\{1/n,...,1/n\}$. However, we stress that to select the best paths $P_{i\rightarrow j}$ or $P_{j\rightarrow i}$ will always consider $w=\{0,...,1\}$. Naturally, we could also consider different combinations of the weights according to the different meanings we aim to give to our application. In other words, we apply the FOWA operator twice when computing betweenness centrality indexes.
As for the fuzzy in-closeness centrality index, also in this case we can consider an alternative measure of the fuzzy out-closeness centrality index as follows:
$\widetilde{C^{(2)}}^{out}_{v_{i},s}=\frac{1}{\widetilde{Z}_{v_{i},s}}$ (38)
In this case, we get an index more similar to the classical non-fuzzy in-closeness centrality index. The interpretation is contrary to 36; the lower the index, the higher the fuzzy in-closeness centrality index. Therefore, in the following, we focus on Equation 36 because the interpretation is quite immediate; indeed, the higher the index in Equation 36, the higher the strength of the fuzzy outcoming edges of a node (given the number of steps to consider the best path), the higher the fuzzy out-closeness centrality of that node.
@@ 6.3.3 Total fuzzy closeness centrality
As we have seen for fuzzy degree centrality, we can also sum up the two different components of closeness centrality in this case. Therefore, we can define the total fuzzy closeness centrality index as follows:
@@ 7 An application of fuzzy social network analysis to collaboration networks in University Departments
For this reason, this study starts from the basic idea of collecting data using the intuition of previous research in a different context. Notably, Calcagnì and Lombardi [10] proposed a new method to collect qualitative data using fuzzy set theory: the so-called Dynamic Fuzzy Rating Tracker (DYFRAT). The authors developed an applet to collect directly triangular fuzzy numbers data without the fuzzy inference process [10]. Their work used mouse movements to model human rating evaluations from a fuzzy-set perspective, capturing the imprecision in point answers. In particular, DYFRAT captures the fuzziness of human ratings by modelling some real-time biometric events that occur during the cognitive rating process the noisy and dynamic x-y trajectory is initially mapped into a set of polar objects (distances and angles). Next, a linear histogram model of angles of movements is built, which stores the most important spatial features of the computer mouse trajectory, such as location, directions, and amplitudes. Finally, the histogram model is used to determine a quantification of spatial events involved in the original movement path. DYFRACT is not available, and thus, in this paper, following the intuition by [10], we built an interactive fuzzy questionnaire to collect fuzzy numbers data to quantify and measure the collaboration relations between professors in the Economic Studies Department of the University “G. d’Annunzio” in Italy.
@@ 8 Discussion e conclusions
In the literature on social network analysis, the links are binary, i.e. present or absent. A possible extension of this approach, widely used in the literature, considers the so-called weighted networks. In the latter, the link between the different nodes of a network is generally expressed with a number in the range $[0,1]$. This approach of considering connections with a weight does not align with fuzzy logic’s philosophy. Fuzzy logic embraces the coexistence of opposites and, for this reason, typically conveys information about vagueness. Attempting to translate an inherently imprecise term (for example, being very good friends) into a single scalar (a number between zero and one) is a strategy that sacrifices information for simplicity.
@@
This work considers the possibility that a vague measure can express the link between different actors in a social network. This approximate measure could be due to an attribute of human language that is often inaccurate by nature; for example, the inaccuracy inherent in concepts such as ”being friends,” ”being little friends,” ”being very close friends,” ”being close collaborators,” or ”having a poor collaboration.” In all these cases, we should have imprecise edges to preserve information about the vagueness of these links. In other cases, as in the application proposed in this work, we can also experience another kind of uncertainty due to indecision in choosing the final qualitative answer to a questionnaire, for example, by checking the mouse’s movements before providing the definitive answer.
@@
The translation of these sources of vagueness has been addressed using triangular fuzzy numbers. This approach extends classic social networks to scenarios where actors are fully connected through nuanced ties, represented by functions. Consequently, this paper introduces a novel method for analyzing social networks, where the connections between actors are expressed through fuzzy numbers. In other words, the links between nodes are represented by functions rather than scalars, as in traditional social network theory.
@@
In light of this new context, this study presented an extension of the traditional centrality indices to the case where the network consists of shaded edges. This research introduces several new measures that consider the imprecision of the links in calculating the centrality measures and, in particular, often provides an original interpretation of these new measures. In the application part of this work, we presented a case of a collaboration network within a university department. The goal was to measure individual collaboration through questionnaires to understand how much each pair of professors or researchers collaborates. The basic idea of this approach is to try to understand the links between department members that go beyond simple indicators such as the number of publications in common or other parameters that can be measured through scalar measures. Indeed, in a department or even in another type of network, there are many cases in which the friendship or, more generally, the collaboration between two members goes beyond the number of tangible products resulting from this collaboration. Strong cooperation between individuals can also exist simply because they have projects in common, exchange ideas, exchange didactic material, go out together outside the working environment, are relatives, and help each other with research questions even if they are not published together, have interests in common outside the university for example for professional activities, or different types of interests in common. In all these cases, considering objective parameters such as the number of joint publications would be highly limiting, and the information would need more effective collaboration between subjects. In any case, since many of these concepts we have listed are by their nature vague, the relationship between individuals and, therefore also, their centrality in a more complex network must be measured through suitable methods that consider the imprecision of human language as well as the uncertainty connected with their way of seeing things. For the latter reason, the social network we propose in this research is rarely symmetrical. In fact, in most cases, an individual’s opinion regarding his relationship with another individual is rarely the same in the opposite direction. For this reason, we have focused on the case of direct social networks in this study. Even if highly improbable, we could have as an extreme case that every link between two nodes is equal in both directions; naturally, as the size of the network or the number of nodes involved increases, the probability that this extreme case occurs tends to zero.
@@
An exciting research component is the possibility of choosing a system of different weights to calculate the proposed centrality indices. Indeed, in some specific contexts and depending on the research objective, it could be interesting to consider other systems of weights to give different importance to solid bonds rather than weak bonds or vice versa. In the theoretical part of the research, we introduced generalized measures to allow the calculation of different centrality measures based on different possible systems of weights. In the application part, there were no particular contextual requirements to think of weights other than those all equal to each other, so we focused on a system of equal weights.
@@
In attributing weights to calculate the centrality indices, choosing the method to create a ranking of fuzzy numbers also assumes great importance. Different methods of sorting fuzzy numbers could lead to different results. This work focused on the centre of gravity method, but our approach can be extended to other procedures. The research focused on triangular fuzzy numbers, but of course, a possible exciting extension would be to consider different membership functions to represent the links existing in a social network. Other interesting future developments could be creating an R package to implement all the functions we have made in this research and developing other indices proposed in the literature on classical social network analysis.
@@
Our approach’s main advantage is to analyse networks of actors in which the relationships are not precise because they are defined by heterogeneous components or simply because they are collected through questionnaires whose possible answers are imprecise or simply attributes of human language. The possible applications include analysing the network of friendships on social networks such as Facebook, Instagram, YouTube, etc. Limiting oneself to simple friendships or followed contacts could be more informative. Actually, being friends with a person on Facebook can have a very nuanced meaning, depending on many other aspects that could be considered. Therefore, considering a relationship between two individuals based on a binary outcome or a simple weight between zero and one would be a complete loss of information because it does not consider the variability or, even better, the so-called fuzziness inherent in the relationship. In this research, we presented an application of the data collected within university departments, but this approach can be extended to multiple real contexts. For example, we could consider workplaces or all networks of individuals involved in sports or government teams. Consequently, the practical applications are numerous and of considerable interest.
@@
No competing interest is declared.
@@ Acknowledgments
The authors thank Dr. Simone Di Nardo for creating the data collection platform and professors Marta Di Nicola and Francesca Scozzari for supporting Dr. Annamaria Porreca in her doctoral path, during which she collected data and worked on the project. We also thank the members of the Department of Economics of the University of Pescara-Chieti for responding to the questionnaire and allowing the study to be carried out.

### v2

@@ 4.3.1 Fuzzy in-closeness centrality
where $\widetilde{A}_{ij}$ is the set of the best fuzzy edges to connect $i$ and $j$ ranked in decreasing order, $s$ indicates the maximum number of steps considered evaluating the best path, and $w_{j}$ a vector of weights of the FTOWA operator.
The fuzzy in-closeness centrality is defined as follows:
@@
Thus, differently from Equation 25 that is a relative frequency (i.e. a crisp number), $\widetilde{C}^{in}_{v_{i},s}$ is a centrality index that incorporates the fuzziness of the edges (i.e. a fuzzy number).
Another possible approach to define fuzzy in-closeness centrality, in a way that is more similar to the classical index, would be to consider $\widetilde{C^{(2)}}^{in}_{v_{i},s}=\frac{1}{\widetilde{M}_{v_{i},s}}$. However, in the following, we refer to Equation 28 because our basic idea is to look for an immediate representation of the in-closeness concept rather than to identically reproduce the classical indexes using fuzzy numbers. In fact, Equation 28 is easily interpretable, because provides a measure of how a node is central in considering the fuzzy incoming edges. However, Equation 28, differently from 16 takes into account the concept of the best path and maximum number of steps to reach another node.
@@ 4.3.2 Fuzzy out-closeness centrality
Following the same reasoning, we can define the fuzzy out-closeness centrality index of the node $v_{i}$ as follows:
@@
where $\widetilde{Z}_{v_{i},s}$ is the average (fuzzy number) of the best paths from nodes $j$ to $i$ in decreasing order. Thus,
@@
is computed by applying the FTOWA operator to fuzzy numbers representing the edges from the generic nodes $j$ and the node $i$, for which we are computing the index.
As before, the vector of weights can be of different types. From now on, we will compute the average fuzzy number $\widetilde{M}_{v_{i},s}$ or $\widetilde{Z}_{v_{i},s}$ using $w=\{1/n,...,1/n\}$. However, we stress that to select the best paths $P_{i\rightarrow j}$ or $P_{j\rightarrow i}$, we will always consider $w=\{0,...,1\}$. Naturally, we could also consider different combinations of the weights according to the different purposes. As for the fuzzy in-closeness centrality index, also in this case we can consider an alternative measure of the fuzzy out-closeness centrality index as $\widetilde{C^{(2)}}^{out}_{v_{i},s}=\frac{1}{\widetilde{Z}_{v_{i},s}}$. This is more similar to the classical non-fuzzy in-closeness centrality index. The interpretation is contrary to 29; the lower the index, the higher the fuzzy in-closeness centrality index. Therefore, in the following, we focus on Equation 29 because the interpretation is quite immediate; indeed, the higher the index in Equation 29, the higher the strength of the fuzzy outcoming edges of a node (given the number of steps to consider the best path), the higher the fuzzy out-closeness centrality of that node.
@@ 4.3.3 Total fuzzy closeness centrality
As we have seen for fuzzy degree centrality, we can also sum up the two different components of closeness centrality. Therefore, we can define the total fuzzy closeness centrality index as follows:
@@ 5 An application of fuzzy social network analysis to collaboration networks in University Departments
For this reason, this study starts from the basic idea of collecting data using the intuition of previous research in a different context. Notably, Calcagnì and Lombardi [2014] proposed a new method to collect qualitative data using fuzzy set theory: the so-called Dynamic Fuzzy Rating Tracker (DYFRAT). The authors developed an applet to collect directly triangular fuzzy numbers data without the fuzzy inference process [Calcagnì and Lombardi, 2014]. Their work used mouse movements to model human rating evaluations from a fuzzy-set perspective, capturing the imprecision in point answers. In particular, DYFRAT captures the fuzziness of human ratings by modelling some real-time biometric events that occur during the cognitive rating process the noisy and dynamic x-y trajectory is initially mapped into a set of polar objects (distances and angles). Next, a linear histogram model of angles of movements is built, which stores the most important spatial features of the computer mouse trajectory, such as location, directions, and amplitudes. Finally, the histogram model is used to determine a quantification of spatial events involved in the original movement path. DYFRACT is not available, and thus, in this paper, following the intuition by [Calcagnì and Lombardi, 2014], we built an interactive fuzzy questionnaire to collect fuzzy numbers data to quantify and measure the collaboration relations between professors in the Economic Studies Department of the University “G. d’Annunzio” in Italy.
@@ 6 Discussion e conclusions
In the literature on social network analysis, the links are binary, i.e. present or absent. A possible extension of this approach, widely used in the literature, considers the so-called weighted networks. In the latter, the link between the different nodes of a network is generally expressed with a number in the range $[0,1]$. This approach of considering connections with a weight does not align with fuzzy logic’s philosophy. Fuzzy logic embraces the coexistence of opposites and, for this reason, typically conveys information about vagueness. Attempting to translate an inherently imprecise human term into a scalar sacrifices information for simplicity.
@@
This work considers the possibility that a vague measure can express the link between different actors in a social network. This approximate measure could be due to an attribute of human language that is inaccurate by nature; for example, the inaccuracy inherent in concepts such as being friends, being little friends, being very close friends, being close collaborators, or having a poor collaboration, etc. In all these cases, we should have imprecise edges to preserve information about the vagueness of these links. In other cases, as in the application proposed in this work, we can also experience another kind of uncertainty due to indecision in choosing the final answer to a questionnaire composed by an ordinal variable, for example, by checking the mouse’s movements before providing the definitive answer. Trying to preserve information about the vagueness of relationships within centrality indices, although the concepts of variability and vagueness are quite different, has something in common with the importance in statistics of preserving information about the variability of the distribution; in other words, it provides us with much richer information than a simple crisp link.
@@
The translation of these sources of vagueness has been addressed using triangular fuzzy numbers. This approach extends classic social networks to scenarios where actors are fully connected through nuanced ties, represented by functions. Consequently, this paper introduces a novel method for analyzing social networks, where the connections between actors are expressed through fuzzy numbers. In other words, the links between nodes are represented by functions rather than scalars, as in traditional social network theory. In light of this new context, this study presented an extension of the traditional centrality indices to the case where the network consists of shaded edges. This research introduces several new measures that consider the imprecision of the links in calculating the centrality measures and, in particular, often provides an original interpretation of these new measures.
@@
In the application part of this work, we presented a case of a collaboration network within a university department. The goal was to measure individual collaboration through questionnaires to understand how much each pair of professors or researchers collaborates. The basic idea of this approach is to try to understand the links between department members that go beyond simple indicators such as the number of publications in common or other parameters that can be measured through scalar measures. Indeed, in a department or even in another type of network, there are many cases in which the friendship or, more generally, the collaboration between two members goes beyond the number of tangible products resulting from this collaboration. Strong cooperation between individuals can also exist because they have projects in common, exchange ideas, exchange didactic material, socialize together outside of the workplace, are relatives, and help each other with research questions even if they are not published together, have interests in common outside the university for example for professional activities, or different types of interests in common. In all these cases, considering objective parameters such as the number of joint publications would be highly limiting. In any case, since many of these concepts we have listed are by their nature vague, the relationship between individuals and, therefore also, their centrality in a more complex network must be measured through suitable methods that consider the imprecision of human language as well as the uncertainty connected with human relationships. For the latter reason, the social network we propose in this research is rarely symmetrical. In fact, in most cases, an individual’s opinion regarding his/her relationship with another individual is rarely the same in the opposite direction. For this reason, we have focused on the case of direct social networks.
@@
An exciting research component is the possibility of choosing a system of different weights to calculate the proposed centrality indices. Indeed, in some specific contexts and depending on the research objective, it could be interesting to consider other systems of weights. In the theoretical part of the research, we introduced generalized measures to allow the calculation of different centrality measures based on different possible systems of weights. In the application part, there were no particular contextual requirements to think of weights other than those all equal to each other, so we focused on a system of equal weights.
@@
In attributing weights to calculate the centrality indices, choosing the method to create a ranking of fuzzy numbers also assumes great importance. Different methods of sorting fuzzy numbers could lead to different results. This work focused on the centre of gravity method, but our approach can be extended to other procedures. The research focused on triangular fuzzy numbers, but of course, a possible exciting extension would be to consider different membership functions to represent the links existing in a social network.
@@
Our approach’s main advantage is to analyse networks of actors in which the relationships are not precise because they are defined by heterogeneous components or simply because they are collected through questionnaires whose possible answers are imprecise or simply attributes of human language. In this research, we presented an application of the data collected within university departments, but this approach can be extended to multiple real contexts. For example, we could consider other workplaces or all networks of individuals involved in sports or government teams. Consequently, the practical applications are numerous and of considerable interest.
@@
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
The authors thank Dr. Simone Di Nardo for contributing to the data collection platform and professors Marta Di Nicola and Francesca Scozzari for supporting Dr. Annamaria Porreca in her doctoral path, during which she collected data and worked on the project. We also thank the members of the Department of Economics of the University of Pescara-Chieti for responding to the questionnaire and allowing the study to be carried out.